# Sharp Bounds for Multiple Models in Matrix Completion

### v2

@@ 3 Conclusion
In this paper, we considered three popular matrix completion settings: (1) the unknown matrix is of low rank and the noise is sub-Gaussian; (2) the unknown matrix is of low rank and the noise is heavy tailed; (3) the unknown matrix is a low rank matrix corrupted by a sparse matrix in the entrywise or columnwise sense, and the noise is sub-Gaussian. We revisited five existing estimators in these settings, and developed sharper upper bounds to establish their minimax rate optimality. The key technical contribution lies in employing powerful matrix concentration inequalities introduced in Brailovskaya and van Handel [2024] to successfully remove the dimension factor $\log d$ that has remained in the upper bounds prior to our work. The $\log d$ factor problem not only exists for the fives estimators studied in this paper, but is ubiquitous throughout the literature. We believe that many other existing results on matrix completion can be possibly improved by applying these new advanced inequalities.
@@ 4 Proof
4.1 Proof of sharp bounds for spectral norms
In this section, we present advanced inequalities introduced in [Brailovskaya and van Handel, 2024] and how to use them. Let $Q=\sum_{i=1}^{n}Q_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ be a random matrix where $Q_{i}$ are independent random matrices with mean zero. We define
$\gamma(Q)$ $=\left(\max(\|EQQ^{T}\|,\|EQ^{T}Q\|)\right)^{1/2},$
$\gamma_{*}(Q)$ $=\sup_{\|y\|_{2}=\|z\|_{2}=1}\left(E(y^{T}Qz)^{2}\right)^{1/2},$
$g(Q)$ $=\|\operatorname{Cov}(Q)\|^{1/2},$
$R(Q)$ $=\max_{1\leq i\leq n}\|Q_{i}\|,$
Theorem 11 (Corrollary 2.17 in [Brailovskaya and van Handel, 2024]).
With the above notations, there exist a universal constant $C$ such that
$P(\|Q\|\geq 2\gamma(Q)+C(g(Q)^{1/2}\gamma(Q)^{1/2}\log^{3/4}d+\gamma_{*}(Q)t^{1/2}+R(Q)^{1/3}\gamma(Q)^{2/3}t^{2/3}+R(Q)t))\leq de^{-t}$ (20)
and a universal constant $C$ such that
The matrix inequalities in [Brailovskaya and van Handel, 2024] are formulated for $Q_{i}\in\mathbb{R}^{d\times d}$ (not necessarily self-adjoint), and can be extended to rectangular matrices in the above form. See Remark 2.1 in [Brailovskaya and van Handel, 2024]. Applying the above inequalities, we have the following results.
@@ 4.1 Proof of sharp bounds for spectral norms
Assume the sampling matrices $X_{i}$ satisfy Assumption 2 and the noise random variables $\xi_{i}$ satisfy Assumption 1. The following inequality holds with probability at least $1-\frac{1}{d}$,
given $n\geq m\log^{6}d\cdot\log^{2}n$. Here $C$ is a constant dependent on $L_{1}\sim L_{3}$. Assume $X_{i}$ satisfy Assumption 2 , $\xi_{i}$ satisfy Assumption 3 and are symmetric. Taking $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$, the following inequality holds with probability at least $1-\frac{1}{d}$,
@@
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\right\|\leq C\max(\sigma,a)\sqrt{\dfrac{1}{nm}}.$ (23)
where $C$ is dependent on $L_{2}$ and $L_{3}$.

### v3

@@ 3 Conclusion
In this paper, we considered three common matrix completion settings: (1) the unknown matrix is of low rank and the noise is heavy tailed; (2) the unknown matrix is of low rank and the noise is sub-Gaussian with known variance; (3) the unknown matrix is of low rank and the noise is sub-Gaussian with unknown variance. We revisited three popular estimators in these settings and developed sharper upper bounds to establish their minimax rate optimality.
A key technical contribution lies in employing powerful matrix concentration inequalities introduced in Brailovskaya and van Handel [2024] to successfully remove the dimension factor $\log d$ that has remained in the upper bounds prior to our work. As previously noted, the $\log d$ factor is ubiquitous in matrix completion problems under the sampling with replacement setting. Our sharp analyses (e.g., Lemma 14 establishes dimension-free results for a variety of random matrices) can be applied or adapted to improve other existing results on matrix completion [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp et al., 2017, Chen et al., 2024].
@@ 4 Proof
Throughout this section, we denote
$L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle).$
The derivative of $l_{\tau}(\cdot)$ is denoted by $\phi_{\tau}(\cdot)$, where
$\phi_{\tau}(x)=\begin{cases}x,|x|\leq\tau,\\ \operatorname{sign}(x)\tau,|x|>\tau.\end{cases}$
In this section, $c_{0}\sim c_{4}$ are specific constants whose values do not change at each occurrence.
4.1 Proof of Theorem 1
We begin with the following general result.
@@ 4.1 Proof of Theorem 1
Let Assumptions 1 and 3 hold. Suppose $n\geq 160m\log^{4}d$ and $\tau=\frac{\max(\sigma,a)}{\log^{2}d}\sqrt{\frac{n}{m}}$. If we take
$\lambda\geq 3\left\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|,$
then the following bound holds with probability at least $1-\frac{1}{d}$,
$\frac{\|\hat{A}_{H}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\mu^{2}rm_{1}m_{2}(\lambda^{2}+a^{2}\mathbb{E}^{2}\|\mathcal{R}\|)+\frac{\mu^{2}\log d}{n}.$
Here $\mathcal{R}$ is defined as $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}},$ where $\chi_{i}=\{|\xi_{i}|\leq\tau/2\}$.
Proof of Theorem 1.
Recall $\tau=\frac{\max(a,\sigma)}{\log^{2}d}\sqrt{\frac{n}{m}}$. By (30), we obtain with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}\right\|\leq C_{1}\max(\sigma,a)\sqrt{\dfrac{1}{nm}},$
where $C_{1}$ depends on $L_{2}\sim L_{3}$. With this specific $C_{1}$, one can take $\lambda=3C_{1}\max(\sigma,a)\sqrt{\frac{1}{nm}}$ to apply Lemma 7. By (32), we obtain
@@
$\mathbb{E}\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\textbf{1}_{\chi_{i}}\right\|\leq C_{2}\sqrt{\frac{1}{nm}},$ (11)
where $C_{2}$ depends on $L_{2}\sim L_{3}$. Substituting this bound and the above choice of $\lambda$ into the inequality from Lemma 7, we complete the proof of Theorem 2. ∎
Remark.
Without the condition that $\xi_{i}$ are symmetric, $\nabla L(A_{0})$ may not have mean 0. In this case, it is necessary to control the spectral norm of its expectation. Specifically, we write:
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\tau}(\xi_{i})X_{i}=\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(\xi_{i})-\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}])X_{i}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}]X_{i}.$
Denote $E[\phi_{\tau}(X_{i})|X_{i}]$ by $u_{i}$ (we omit the dependence on $\tau$ to ease the notation). While (30) still applies to $\frac{1}{n}\sum_{i=1}^{n}(\phi_{\tau}(\xi_{i})-u_{i})X_{i}$ given $\tau=\max(\sigma,a)\sqrt{\frac{n}{m\log^{4}d}}$, we need to estimate
$\left\|\frac{1}{n}\sum_{i=1}^{n}u_{i}X_{i}\right\|$ $=\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[u_{i}X_{i}]\right\|$
$\leq\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|+\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[u_{i}X_{i}]\right\|.$
Assume a random variable $Y$ has 0 mean and finite moment of order $2+\kappa$ with $\mathbb{E}^{\frac{1}{2+\kappa}}[|Y|^{2+\kappa}]\leq\sigma$ where $\kappa\geq 0$. By Cauchy-Schwarz inequality, for a positive number $\tau$,
$|\mathbb{E}\phi_{\tau}(Y)|$ $=|\mathbb{E}[Y-\phi_{\tau}(Y)]|$
$=|\mathbb{E}[(Y-\phi_{\tau}(Y)\textbf{1}(|Y|>\tau)]$
$\leq\mathbb{E}^{1/p}[|Y|^{p}]\mathbb{P}^{1/q}(Y>\tau)$
where $11/p+1/q=$ 1. Note $\mathbb{P}(Y>\tau)\leq\mathbb{E}[|Y|^{2+\kappa}]/\tau^{2+\kappa}$. Take $1/q=1-1/(2+\kappa)$. Then, we obtain
$|\mathbb{E}[(Y-\phi_{\tau}(Y))\textbf{1}(|Y|>\tau)]|\leq\mathbb{E}^{\frac{1}{2+\kappa}}[|Y|^{2+\kappa}](\mathbb{E}^{\frac{1+\kappa}{2+\kappa}}[|Y|^{2+\kappa}]\tau^{-(1+\kappa)}).$
Let $\tau=\max(\sigma,a)\sqrt{\frac{n}{m\log^{4}d}}$. By the above inequality, we obtain $|u_{i}|\leq\max(\sigma,a)(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$. Assuming $P_{jk}=\frac{1}{m_{1}m_{2}}$, we apply Theorem 7.3.1 in Tropp et al. [2015] to $\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])$. Since $\frac{1}{n}\|u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}]\|$ is bounded by $\max(\sigma,a)\frac{2}{n}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$, the following holds with probability at least $1-\frac{1}{d}$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|\lesssim\max(\sigma,a)\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2}\sqrt{\frac{\log d}{nm}}+\max(\sigma,a)\frac{\log d}{n}\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2},$
where we use easy estimates $\|\mathbb{E}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])^{T}\|\lesssim\max(\sigma,a)\frac{L_{2}}{m}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$ and $\|\mathbb{E}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])^{T}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\|\lesssim\max(\sigma,a)\frac{L_{2}}{m}(\frac{m\log^{4}d}{n})^{(1+\kappa)/2}$. Given $n\geq m\log^{4+1/(1+\kappa)}d$, we have
$\left\|\dfrac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])\right\|\lesssim\sigma\sqrt{\frac{1}{nm}}.$
Denote $u_{i}(j,k)=\mathbb{E}[\phi_{\tau}(\xi_{i})|X_{i}=e_{j}e^{T}_{k}]$. Note
$\|\mathbb{E}[u_{i}X_{i}]\|$ $=\sup_{\|z\|_{2}=1,\|y\|_{2}=1}z^{T}\mathbb{E}[u_{i}X_{i}]y$
$\leq\left(\sum_{j,k}P^{2}_{jk}\right)^{1/2}\left(\sum_{j,k}z_{j}^{2}y_{k}^{2}u^{2}_{i}(j,k)\right)^{1/2}$
$\leq\max(\sigma,a)\sqrt{\frac{1}{m_{1}m_{2}}}\left(\frac{m\log^{4}d}{n}\right)^{(1+\kappa)/2}.$
Given $Mn^{\kappa}\geq m\log^{4}d$, the RHS is less than $\max(\sigma,a)\sqrt{\frac{1}{nm}}$. An alternative condition is $n\geq\log^{4/\kappa}d$. When the noise only has second moment, namely $\kappa=0$, we can derive a $\log d$-free bound given $M\geq m\log^{4}d$ and $n\geq m\log^{5}d$.
For non-symmetric noise, a $\log d$-free bound may be obtained under a better condition (i.e. $M\geq m\log^{\beta}d$ where $54\leq\beta<$ 5) by applying Theorem 13 to $\frac{1}{n}\sum_{i=1}^{n}(u_{i}X_{i}-\mathbb{E}[u_{i}X_{i}])$. We do not make further efforts here since it cannot remove the condition entirely. This technical requirement will possibly be not necessary with future refinement of the sharp concentration inequalities in Brailovskaya and van Handel [2024].
Before presenting the poof of Lemma 7, we state a technical lemma.